OpenStudy (anonymous):

what is the min or max of y=x^2-10x+25

OpenStudy (solomonzelman):

I think they are asking for the vertex because they say the minimum OR maximum can you find the vertex?

OpenStudy (anonymous):

I dont understand how to

OpenStudy (solomonzelman):

I stink at explaining, but I have this http://www.algebra.com/algebra/homework/Rational-functions/Rational-functions.faq.question.501396.html you can also put in your equation ion the bottom and the solution will be given....

OpenStudy (anonymous):

Given equation is an expansion of (x - 5)^2 with solution 5 so max value is 5

OpenStudy (anonymous):

can you explain how you got that

OpenStudy (anonymous):

(x -5)^2 = x^2 - 10x + 25 (a - b)^2 = a^2 - 2ab + b^2

OpenStudy (anonymous):

x - 5 = 0 and x - 5 = 0 x = 5 Which is the max value for the equation

OpenStudy (anonymous):

so what would be the vertex

OpenStudy (anonymous):

Min or max value of function is given by c = -b^2/4a c = -100/4 = -25 which means the given function has minimum value of -25